One machine in a print shop can produce a cetain # of pages twice as fast as another machine. Operating together these machines can produce this number of pages in 8 minutes. How long would it take the slower machine working alone to produce this number of pages?
______minutes.
Thanx

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLES AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
-----------------------------------------------------------------
Mixture_Word_Problems/30424: Dennis can eat a pie in 6 min. David can eat a pie in 12 min. Robert can eat a pie in 14.639145073927682623334797247841 min. If they share the same pie, how lomg will it take the three guys to scarf down the whole thing. (Hint: First figure out how many pies each guy can eat in an hour)
1 solutions
Answer 17065 by venugopalramana(1167)on 2006-03-15 22:41:46 
Dennis can eat a pie in 6 min. David can eat a pie in 12 min. Robert can eat a pie in 14.639145073927682623334797247841 min. If they share the same pie, how lomg will it take the three guys to scarf down the whole thing. (Hint: First figure out how many pies each guy can eat in an hour)
DENIS CAN EAT IN 1 MINUTE....1/6 =0.166666667 PIE
DAVID CAN EAT IN 1 MINUTE....1/12 =0.083333333 PIE
ROBERT CAN EAT IN 1 MINUTE...1/14.63914507=0.06831 PIE
HENCE THE 3 TOGETHER CAN EAT IB 1 MINUTE=0.31831 PIE
HENCE THEY TAKE ...1/0.31831 = 3.14159153 MINUTES TO EAT THE PIE.
Rate-of-work-word-problems/29587: need help plz...heres the question "There are three friends named Allan, Bobby and Charlie. The three friends want to know their individual rate in finishing a job. Allan and Bobby can finish the job in 42 days, Bobby and Charlie can finish the job in 31 days, and Allan and Charlie can finish the job in 20 days. Solve the rate of each individual."
1 solutions
Answer 16715 by venugopalramana(1167)on 2006-03-12 11:04:12 
need help plz...heres the question "There are three friends named Allan, Bobby and Charlie. The three friends want to know their individual rate in finishing a job. Allan and Bobby can finish the job in 42 days, Bobby and Charlie can finish the job in 31 days, and Allan and Charlie can finish the job in 20 days. Solve the rate of each individual."
LET ALLAN TAKE A DAYS TO DO THE JOB ALONE
HENCE ALLAN ALONE CAN DO IN 1 DAY 1/A JOB
LET BOBBY TAKE B DAYS TO DO THE JOB ALONE
HENCE BOBBY ALONE CAN DO IN 1 DAY 1/B JOB
LET CHARLIE TAKE C DAYS TO DO THE JOB ALONE
HENCE CHARLIE ALONE CAN DO IN 1 DAY 1/C JOB
FROM ABOVE WE GET ...
ALLAN AND BOBBY CAN DO IN 1 DAY ..1/A +1/B =(A+B)/AB JOB..
SO DAYS THEY TAKE TO COMPLETE THE JOB =AB/(A+B)=42..OR..1/A+1/B=1/42..........I BOBBY AND CHARLIE CAN DO IN 1 DAY ..1/B +1/C =(B+C)/BC JOB..
SO DAYS THEY TAKE TO COMPLETE THE JOB =BC/(B+C)=31..OR..1/B+1/C=1/31.........II
ALLAN AND CHARLIE CAN DO IN 1 DAY ..1/A +1/C =(A+C)/AC JOB..
SO DAYS THEY TAKE TO COMPLETE THE JOB =AC/(A+C)=20..OR..1/A+1/C=1/20........III
EQNI+EQNII+EQNIII GIVES...
2{(1/A)+(1/B)+(1/C)}=(1/42)+(1/31)+(1/20)
(1/A)+(1/B)+(1/C)=(1/2)*{(1/42)+(1/31)+(1/20)}.........IV
EQN.IV-EQN.I GIVES
1/C= (1/2)*{(1/42)+(1/31)+(1/20)}-1/42....OR.....C=34.2 DAYS
EQN.IV-EQN.II GIVES.......
1/A=(1/2)*{(1/42)+(1/31)+(1/20)}- 1/31...OR......A=48.1 DAYS
EQN.IV-EQN.III GIVES....
1/B=(1/2)*{(1/42)+(1/31)+(1/20)}- 1/20....OR.....B=329.6 DAYS
---------------------------------------------------------------------------
Pipe A can fill a tank in
8 hours. When pipe A and pipe B are used, the tank is
filled in 5 hours. How long, in hours, would it take
pipe B alone to fill the tank?
This is not from a book; it is from a handout.
Thank you for any help. I appreciate it.
1 solutions
Answer 22576 by venugopalramana(1834)on
2006-05-08 05:18:51 
SEE THE FOLLOWING EXAMPLE AND TRY
Could you please help me
with this word problem. I'm really having a hard time
with it: Pipes A and B can fill a storage tank in 8
hours and 12 hours. With the tank empty, pipe A was
turned on at noon, and then pipe B was turned on at
1:30 P.M. At what time was the tank full?
Thank you.
1 solutions
Answer 13931 by venugopalramana(1619)on
2006-01-30 05:52:42 
Pipes A and B can fill a storage tank in 8 hours and
12 hours.
THAT IS PIPE A CAN FILL IN 1 HR.1/8 TANK AND
PIPE B CAN FILL IN 1 HOUR 1/12 TANK
With the tank empty, pipe A was turned on at noon, and
then pipe B was turned on at 1:30 P.M.
HENCE IN 1=30 PM - 12 = 00 NOON =1.5 HRS PIPE A FILLS
1.5*1/8 TANK=1.5/8 TANK
REST TO BE FILLED =1-1.5/8=6.5/8 TANK
WHEN A AND B ARE OPEN IN 1 HOUR THEY FILL 1/8 +1/12
TANK = 5/24 TANK
HENCE TO FILL 6.5/8 TANK THEY NEED
{(6.5/8)/(5/24)}=(24*6.5)/(8*5)=3.9 HRS.=3 HRS 54 MTS
HENCE THE TANK WILL BE FULL AT
1=30.....+.....3=54.......5=24 PM
At what time was the tank full?
The inlet pipe of an oil tank can fill the
tank in 1 hour, 30 minutes. The outlet pipe can empty
the tank in 1 hour. How long does it take to empty the
tank if both pipes are open?inlet pipe fills 1 tank in
1hr 30 mts =1.5=3/2 hrs.or inlet pipe fills
...1/(3/2)=2/3 tank in 1 hr.out let pipe empties 1
tank in 1 hr....so assuming the tank is full at the
start water to be emptied =1 tank and in 1 hr inlet
fills 2/3 tank and out let empties 1 tank ..so net out
flow is 1-2/3=(3-2)/3=1/3 tank in 1 hour.so 1/3 tank
emptying needs .....1 hr then 1 tank emptying needs
......? hrs...= 1*1/(1/3)=3 hrs.